Please consider the rational function: f(x)=8x+1/x^2-x-6
Find all vertical and horizontal asymptotes; give the equation(s) of the lines(s) as your answer.

Answer by lwsshak3(11628)

Please consider the rational function: f(x)=8x+1/x^2-x-6
Find all vertical and horizontal asymptotes; give the equation(s) of the lines(s) as your answer.
**
f(x)=8x+1/x^2-x-6
since degree of denominator is one less than degree of numerator,
horizontal asymptote: y=0, or x-axis
...
Vertical asymptote:
set denominator=0, then solve for x-values which make the function undefined
x^2-x-6=0
(x-3)(x+2)=0
x≠3 or x≠-2
vertical asymptotes: x=3 and x=-2